One-half of Nicole's age two years from now, plus one-third of her age three years ago is twenty. How old is she now?
Answer by stanbon(75887)

One-half of Nicole's age two years from now, plus one-third of her age three years ago is twenty. How old is she now?
----
Equation:
(1/2)(n+2)+(1/3)(n-3) = 20
Multiply thru by 6 to get rid of the fractions:
3(n+2) + 2(n-3) = 120
3n+6 + 2n-6 = 120
5n = 120
n = 24 yrs (Nicole's age now)
==============
Cheers,
Stan H.
=============